Two particles of equal mass move around a circle of radius R under the action of their mutual gravitational attraction.What is the speed v of each particle?

Grade:12

1 Answers

Nirmal Singh.
10 years ago
Both mass will move with equal velocity the mutual gravitational attraction force with be equal to centripetal force

so Gmm / (2R)^2 = mv^2 / R

On solving v = sqrt (Gm / 4R^3)

Regards,
Nirmal Singh
Askiitians Faculty